# Scalable Training of Graph Foundation Models for Atomistic Materials Modeling: A Case Study with HydraGNN

### v2

@@ IV-B I/O Performance for Reading Large Data
Of all the datasets used in this study, the Open Catalyst 2020 dataset is the largest in terms of the number of molecules, the storage size of the dataset, and the number of files across which data is stored. The original dataset consists of over 50,000 files. The dataset was pre-processed into ADIOS and was configured to use just over 50 ADIOS sub-files, which led to a 1000$\times$ reduction in the metadata footprint.
When training begins, HydraGNN reads ADIOS data in parallel on all processes. This read operation is a two-step process in which the root process first obtains the number of graphs (molecules) followed by the size (number of atoms) and the feature metadata for each graph. This information is broadcast to all other processes that implicitly distribute the graphs evenly amongst themselves and concurrently read their assigned graphs from the ADIOS file. This set of operations is performed for all molecule groups - trainset, valset, and testset.
Fig. 3 shows the reading performance of trainset data on Frontier when all processes read their assigned graphs in parallel. We obtain over 8 Terabytes/second for higher node counts and almost 2 Terabytes/second on 128 nodes. The high I/O bandwidth is a characteristic feature of the ADIOS library as it permits multiple processes to read data spread over multiple ADIOS sub-files efficiently. A similar run on Perlmutter was not possible due to filesystem issues encountered on the system during our study.
@@
Figure 4: Read performance for the entire OC2020 dataset on Frontier that includes training, validation, and testing data. We obtain over 120 Gigabytes/second (approximately 35 seconds) for reading 4.3 Terabytes of data.
@@ IV-C1 Weak Scaling
For the weak scaling runs, we configured each GPU to process 3,500 molecules equally. Fig. 5 shows the weak scaling performance on Perlmutter and Frontier as we vary in the number of GPUs used for the training. The reported time represents the average training time per epoch. We conducted experiments with up to 2,048 GPUs on Frontier and 1,024 GPUs on Perlmutter. The limited number of GPUs on Perlmutter was due to constraints on available node hours. We observe that the parallel efficiency of weak scaling experiments drops as we increase the number of GPUs beyond 256 for both Perlmutter and Frontier. This is attributed to increased communication costs as we scale the number of GPUs and the overhead associated with using varying graph sizes.
Figure 5: Weak scaling of HydraGNN multitasking pretraining on a problem of size 3,500 molecules per GPU for (top) Frontier and (bottom) Perlmutter.
Fig.  6 provides a breakdown of the overhead of different components of HydraGNN used in the weak scaling experiments. ‘forward’ and ‘backward’ represent the forward and backward phases of the DL model training, respectively, and ‘dataload’ denotes the cost of obtaining the next batch of data samples from DDStore after a GPU finishes processing its current batch. We notice that ‘dataload’ has a fixed cost, which expectedly becomes more prominent for the small model size and is only a fraction of the runtime as the model size increases. The forward and backward phases show an increase in runtime as we scale up the workflow as synchronization and communication operations become more expensive with increasing GPU counts.
@@ IV-C2 Strong Scaling
For strong scaling runs, we trained HydraGNN on 120 million molecular graphs (approximately 4TB in size) on Frontier and 2 million molecular graphs on Perlmutter for the three model sizes. Fig. 7 shows the scaling results for 512 to 16,384 GPUs on Frontier and from 64 up to 2,048 GPUs on Perlmutter. The reported time is the average training time per epoch, similar to the weak scaling measurements. While the SMALL model’s performance deviates from the optimal linear dotted line after 2,048 GPUs on Frontier, the MEDIUM and LARGE models maintain close to linear scaling up to 16,384 GPUs on Frontier. We notice a similar trend on Perlmutter where we observe near-linear scaling upto 2048 GPUs for all model sizes.
@@
The drop in scaling performance is attributed to load imbalance - an artifact of varying graph sizes. As shown in Fig. 1, we use a diverse dataset where molecule sizes vary by up to 400, and the number of edges in the larger molecules exceeds 12,500. This results in an imbalanced workload among GPUs in each batch, causing some GPUs to finish training before others. As GPUs must synchronize for exchanging model weights, the runtime is dominated by the GPUs that must train on larger graphs. Effectively, this leads to sub-standard utilization of compute resources and posts a challenge towards achieving high-performant, scalable training. Training on large volumes of data can help develop robust models because of the diverse nature of data, but it affects the computational performance as the workload can vary greatly.
Figure 8: The distribution of forward time during the training of three models with respect to the graph size, measured in the number of edges. It illustreate a linear relationship between forward time and graph size.
The EGNN model we used is particularly vulnerable to this problem. The time required for forward calculations in EGNN is directly proportional to the number of edges in the molecules. For datasets with highly variable edge counts between molecules, the likelihood of load imbalance between GPUs increases. Fig. 8 illustrates the time spent on the forward task in the EGNN model with different model sizes. We observe an almost linear relationship between forward time and graph size at each batch (measured by the number of edges). The SMALL models show large variances on both machines, which is expected due to system noise being more apparent with smaller model sizes. Significant performance differences (e.g., the difference between minimum and maximum time) are observed due to the varying graph sizes in our datasets. However, for other tasks (data loading and backward), we do not observe a similar correlation, as they are agnostic of the graph size. Fig. 9 illustrates the average percentage of time spent waiting during three tasks: data loading, forward pass, and backward pass. It highlights a significant waiting period during the forward pass, primarily due to varying graph sizes. This waiting time increases as the disparity in graph sizes among GPUs grows. Other tasks, such as data loading and backward pass, also involve waiting time, but to a much lesser extent.
@@
To quantify the degree of load imbalance between GPUs, we compute the Load Imbalance Factor (LIF). The LIF is determined by the ratio of the maximum load observed on a computing resource to the average load across all resources. We define LIF as
@@
where $T_{max}$ and $T_{avg}$ represent the maximum runtime and the average runtime for training an epoch, respectively, among all computing resources (GPUs in our case). These timings represent the time to perform training (forward and backward calculations) and do not include wait times during synchronization. For a well-balanced workload, LIF approaches 1.0 from above, whereas it increases as the workload imbalance increases. Fig. 10 presents the LIF scores to demonstrate the imbalance among processes. The trend remains consistent: while data loading and backward pass exhibit nearly balanced workloads (with scores close to 1.0), the forward pass shows imbalanced workload characteristics as it deviates from 1.0.
@@ IV-D Scalable HPO
The HPO process is performed using DeepHyper [79], a Python package that provides a common interface for the implementation and study of scalable hyperparameter search methods. DeepHyper has been specifically designed to perform efficient and scalable HPO on integrated extreme scale HPC and leadership class supercomputing facilities, and it thus suits very well for our purpose. Among the various hyperparameter search algorithms implemented in DeepHyper, we used the Centralized Bayesian Optimisation Search [86, 87, 88, 89, 90, 91], previously named as “Asynchronous Model-Based Search” (AMBS) [92]. It follows a manager-workers architecture where the manager runs the Bayesian optimization loop and workers execute parallel evaluations of the black-box function.
@@
We have successfully executed HPO using 8,192 Frontier nodes in parallel. Each HPO trial is associated with an independent ‘srun’ execution of the SLURM scheduler and occupies 256 nodes (i.e., 2,048 AMD Mi250x GCDs) for distributed training using DDP, thereby allowing 32 distinct HydraGNN architectures to be concurrently trained. Concurrent HPO trials are executed asynchronously, and the termination of an HPO trial is immediately followed by the start of a new one on the same set of compute nodes. This extensive scale not only tests the limits of scalability and efficiency in computational resources, but also addresses the challenges associated with the high dimensionality of the hyperparameter space that needs to be explored, ensuring that a proper balance between exploitation and exploration is maintained during the entire HPO execution.
@@
In order to ensure that the HPO process is performed in an energy-efficient way on OLCF-Frontier, we early stop the training of HydraGNN models for each HPO trial after 10 epochs. This number of epochs allows to early stop the HPO trials that are clearly underperforming in a timely manner, without wasteful energy consumption caused by further training epochs that would not likely improve their accuracy, while still ensuring that promising HPO trials are distinguishable and selected for the next computational tasks. This approach results into impactful energy savings. Our use of DeepHyper for asynchronous Bayesian optimization, combined with a strategic deployment of early termination strategies, showcases a significant advancement in the field, optimizing GNN training in ways that have not been documented prior to this.
@@
Fig. 11 reports the validation mean absolute error (MAE) as a function of wall-clock time. The scattered distribution of blue dots (corresponding to values of the validation MAE for different HPO trials) shows that the HPO maintains a good degree of exploration throughout the entire execution. The red solid line indicates the minimum validation MAE obtained at a given time during the HPO run. The fact that the red line progressively lowers as time progresses confirms that new HPO trials progressively selected and evaluate help identify GFM architecture with better accuracy.
@@
Figure 11: Value of validation MAE for different HPO trials as a function of wall-clock time expressed in seconds.
We found that the 10 best performing HydraGNN models identified by HPO are relatively small in size, between 4 and 6 millions parameters. The fact that HPO proposes models with small number of parameters seems to disagree with other results presented in the literature [35] for GFM applied to atomistic materials modeling, where models with increasing numbers of parameters (up to a few billions of parameters) seem to increasingly reach higher accuracy, even when they are trained on smaller volumes of data than what we use for our study. One possible explanation for this disagreement is that smaller models tend to learn faster, thus inducing HPO to favor smaller models when early stopping is applied to each HPO trial.
We also found that all the 10 best HydraGNN models selected by HPO use the PNA as MPNN layer. Among the different types of MPNN lyers tested, the PNA (albeit non-invariant and non-equivariant) was already shown to reach higher accuracy on alloy systems [38, 93, 73]. Since the OC2022 and the OC2022 datasets, obtained by slicing 3D bulk alloy systems in 2D slabs and modelling the interaction with catalysts on the 2D alloy surface, constitute the majority of the training data, the automated HPO analysis performed in this study seems to reconfirm what was already empirically observed at smaller scale in previous studies conducted by the authors.
@@
To characterize the dynamic resource behavior of HPO, the user-space telemetry tool highlighted in Section IV-A was enabled to capture a variety of GPU metrics during a small HPO exercise utilizing 320 GPUs. Fig. 12 highlights the memory subsystem showing individual GPU memory usage on each assigned compute node with a dynamic high water mark peaking at 99% of available memory. The variability of memory utilization across different GPUs is due to the fact that different groups of GPUs (associated with different HPO trials) train HydraGNN models of different size, which affects the amount of GPU memory engaged at different stages of the model training.
@@
Figure 12: GPU HBM memory consumption traces sampled via omnistat telemetry harness (per GCD on each assigned compute node) during HPO exercise executed on OLCF-Frontier.
@@ IV-E Energy Profiling
To quantify the energy usage as a function of different model sizes, three training epochs of the SMALL, MEDIUM, and LARGE model configurations listed in Table III were completed with the omnistat telemetry tool sampling at one second intervals. Measurements consider the entire run duration including I/O for the initial data loading process. Each model executed using 1,024 GPUs on 128 compute nodes which is the minimum node count needed to accommodate memory requirements for the LARGE model configuration. The resulting GPU energy measurements as a function of model size are summarized in Table V. While total energy scales with the execution time, note that GPU utilization (occupancy) also influences the energy consumed. Table V includes mean utilization observed across all 1,024 GPUs. The LARGE case showed the highest GPU utilization–around 89%. The underling power histories used to compute total energy consumed for each model configuration are shown in Fig. 13. From these plots, we see evidence of the underlying training process with three epoch cycles visible in the power response. Furthermore, the increased GPU utilization for the larger models leads to increased GPU power demand with the LARGE model encountering peak power measurements in excess of 520 W (the peak TDP power for the AMD MI250 socket is 560W).
@@
Figure 13: GPU Energy use over time for three models – SMALL (top), MEDIUM (middle), and LARGE (bottom). Each line represents one AMD MI250x.
@@
IV-F Full training of best performing HydraGNN models identified by scalable HPO

### v3

@@ IV-B I/O Performance for Reading Large Data
Of all the datasets used in this study, the Open Catalyst 2020 dataset is the largest in terms of the number of atomistic structures, the storage size of the dataset, and the number of files across which data is stored. The original dataset consists of over 50,000 files. The dataset was pre-processed into ADIOS and was configured to use just over 50 ADIOS sub-files, which led to a 1000$\times$ reduction in the metadata footprint.
When training begins, HydraGNN reads ADIOS data in parallel on all processes. This read operation is a two-step process in which first the root process obtains the number of graphs (atomistic structures) followed by the size (number of atoms) and the feature metadata for each graph. This information is broadcast to all other processes that implicitly distribute the graphs evenly amongst themselves and concurrently read their assigned graphs from the ADIOS file. This set of operations is performed for all atomistic structures groups - training set, validation set, and testing set.
Fig. 7 shows the reading performance of the training data on Frontier when all processes read their assigned graphs in parallel. We obtain over 8 Terabytes/second for higher node counts and almost 2 Terabytes/second on 128 nodes. The high I/O bandwidth is a characteristic feature of the ADIOS library as it permits multiple processes to read data spread over multiple ADIOS sub-files efficiently. A similar run on Perlmutter was not possible due to PFS issues encountered on the system during our study.
@@
Figure 8: Read performance for the entire OC2020 dataset on Frontier that includes training, validation, and testing data. We obtain over 120 Gigabytes/second (approximately 35 seconds) for reading 4.3 Terabytes of data. We will apply techniques to read user metadata in parallel to improve the overall read performance of the complete dataset.
@@ IV-C1 Weak Scaling
For the weak scaling runs, we configured each GPU to process 3,500 atomistic structures equally. Fig. 9 shows the weak scaling performance on Perlmutter and Frontier as we vary the number of GPUs used for the training. The reported time represents the average training time per epoch. We conducted experiments with up to 2,048 GPUs on Frontier and 1,024 GPUs on Perlmutter. The number of GPUs on Perlmutter was due to constraints on available node hours. We observe that the parallel efficiency of weak scaling experiments drops as we increase the number of GPUs beyond 256 for both Perlmutter and Frontier. This is attributed to increased communication costs as we scale the number of GPUs and the overhead associated with using varying graph sizes.
Figure 9: Weak scaling of HydraGNN multitasking pretraining on a problem of size 3,500 atomistic structures per GPU for (top) Frontier and (bottom) Perlmutter.
Fig.  10 provides a breakdown of the overhead of different components of HydraGNN used in the weak scaling experiments. The terms ‘forward’ and ‘backward’ represent the forward and backward phases of the DL model training, respectively, and ‘dataload’ denotes the cost of obtaining the next batch of data samples from DDStore after a GPU finishes processing its current batch. We notice that ‘dataload’ has a fixed cost, which expectedly becomes more prominent for the small model size and is only a fraction of the runtime as the model size increases. The forward and backward phases show an increase in runtime as we scale up the workflow as synchronization and communication operations become more expensive with increasing GPU counts.
@@ IV-C2 Strong Scaling
For strong scaling runs, we trained HydraGNN on a subset of the entire data available, comprising 120 million atomistic graphs (approximately 4TB in size), on Frontier and 2 million atomistic graphs on Perlmutter for the three model sizes. Fig. 11 shows the scaling results for 512 to 16,384 GPUs on Frontier and from 64 up to 2,048 GPUs on Perlmutter. The reported time is the average training time per epoch, similar to the weak scaling measurements. While the SMALL model’s performance deviates from the optimal linear dotted line after 2,048 GPUs on Frontier, the MEDIUM and LARGE models maintain close to linear scaling up to 16,384 GPUs on Frontier. We notice a similar trend on Perlmutter where we observe near-linear scaling up to 2,048 GPUs for all model sizes.
@@
The drop in scaling performance is attributed to load imbalance - an artifact of small graph sizes. As shown in Fig. 1, we use a diverse dataset where graph sizes vary by up to 400 nodes, and the number of edges in the larger graphs exceeds 12,500. This results in an imbalanced workload among GPUs in each batch, causing some GPUs to finish training before others. As GPUs must synchronize for exchanging model weights, the runtime is dominated by the GPUs that must train on larger graphs. Effectively, this leads to sub-standard utilization of compute resources and poses a challenge towards achieving high-performant, scalable training. Training on large volumes of data can help develop robust models because of the diverse nature of data, but it affects the computational performance as the workload can vary greatly.
Figure 12: The distribution of forward time during the training of three models with respect to the graph size, measured in the number of edges. It illustrate a linear relationship between forward time and graph size.
The EGNN model we used is particularly vulnerable to this problem. The time required for forward calculations in EGNN is directly proportional to the number of edges in the graph sample. For datasets with highly variable edge counts between graph samples, the likelihood of load imbalance between GPUs increases. Fig. 12 illustrates the time spent on the forward task in the EGNN model with different model sizes. We observe an almost linear relationship between forward time and graph size at each batch (measured by the number of edges). The SMALL models show large variances on both machines, which is expected due to system noise being more consequential with smaller model sizes. Significant performance differences (e.g., the difference between minimum and maximum time) are observed due to the varying graph sizes in our datasets. However, for other tasks (data loading and backward), we do not observe a similar correlation, as they are agnostic of the graph size. Fig. 13 illustrates the average percentage of time spent waiting during three tasks: data loading, forward pass, and backward pass. It highlights a significant waiting period during the forward pass, primarily due to varying graph sizes. This waiting time increases as the disparity in graph sizes among GPUs grows. Other tasks, such as data loading and backward pass, also involve waiting time, but to a much lesser extent.
@@
To quantify the degree of load imbalance between GPUs, we compute the Load Imbalance Factor (LIF) defined by the ratio
@@
where $T_{max}$ and $T_{avg}$ represent the maximum runtime and the average runtime for training an epoch, respectively, among all computing resources (GPUs in our case). These timings represent the time to perform training (forward and backward calculations) and do not include wait times during synchronization. For a well-balanced workload, LIF approaches 1.0 from above, whereas it increases as the workload imbalance increases. Fig. 14 presents the LIF scores that show the imbalance among processes. The trend remains consistent: while data loading and backward pass exhibit nearly balanced workloads (with scores close to 1.0), the forward pass shows imbalanced workload characteristics as it deviates from 1.0.
@@ IV-D Scalable HPO
The HPO process is performed using DeepHyper [84]. DeepHyper has been specifically designed to perform efficient and scalable HPO on integrated extreme scale HPC and leadership class supercomputing facilities, and it thus suits well our purpose. Among the various hyperparameter search algorithms implemented in DeepHyper, we used the Centralized Bayesian Optimisation Search [93, 94, 95, 96, 97, 98], previously named as “Asynchronous Model-Based Search” (AMBS) [99]. It follows a manager-workers architecture where the manager runs the Bayesian optimization loop and workers execute parallel evaluations of the black-box function.
@@
We ran four consecutive HPO runs of progressively increasing scale, each restarting from the output of the previous one. During each one of the four HPO runs, each HPO trial is associated with an independent ‘srun’ execution of the SLURM scheduler and occupies 128 Frontier nodes (i.e., 1,024 AMD Instinct MI250x GCDs) for distributed training using DDP. Concurrent HPO trials are executed asynchronously, and the termination of an HPO trial is immediately followed by the start of a new one on the same set of compute nodes. Throughout the entire execution of each HPO run, we used the open-source Omnistat data-collection infrastructure highlighted in Section IV-A to simultaneously collect GPU telemetry with measurements saved on a local Lustre PFS. Since each HPO trial is submitted as a separate job step within the global job submission, Omnistat post-processing allows us to isolate power and energy measurements at the job-step level thus providing an aggregate GPU energy consumed estimate for each trial independently. In order to ensure that the HPO process is performed in an energy-efficient way on OLCF-Frontier, we early stop the training of HydraGNN models for each HPO trial after 10 epochs. This number of epochs allows to early stop the HPO trials that are clearly underperforming in a timely manner, without wasteful energy consumption caused by further training epochs that would not likely improve their accuracy, while still ensuring that promising HPO trials are distinguishable and selected for the next computational tasks. This approach results into impactful energy savings. Our use of DeepHyper for asynchronous Bayesian optimization, combined with in-band telemetry monitoring and a strategic deployment of early termination strategies, showcases a significant advancement in the field, optimizing GNN training in ways that have not been documented prior to this work.
The first HPO run used 2,048 Frontier nodes in parallel, thereby allowing 16 distinct HydraGNN architectures to be concurrently trained. This first HPO run allowed us to perform a first uninformed exploration of the hyperparameter space, and construct some guidance for the consecutive HPO runs. The results of the first HPO run have been used as inputs to inform the second HPO run, which used 3,072 Frontier nodes in parallel, thereby allowing 28 distinct HydraGNN architectures to be concurrently trained. The results of the second HPO run have been used as input to inform the third HPO run, which use 4,096 Frontier nodes in parallel, thereby allowing 32 distinct HydraGNN architectures to be concurrently trained. Similarly, the output of the third HPO run has been used to successfully guide the fourth and last HPO run, which used 8,560 Frontier nodes in parallel, thereby allowing 67 distinct HydraGNN architectures to be concurrently trained. This extensive scale of the last HPO run not only tests the limits of scalability and efficiency in computational resources, but also addresses the challenges associated with the high dimensionality of the hyperparameter space that needs to be explored, ensuring that a proper balance between exploitation and exploration is maintained.
@@
Figures 15, 16, 17, and 18 show the validation mean absolute error (MAE) as a function of wall-clock time for each one of the four HPO runs. For each one of the four HPO runs, the scattered distribution of blue dots (corresponding to values of the validation MAE for different HPO trials) shows that the HPO maintains a good degree of exploration throughout the entire execution. The red solid line indicates the minimum validation MAE obtained at a given time during the HPO run. The fact that the red line progressively lowers as time progresses confirms that HPO progressively identifies GFM architectures with better accuracy. Moreover, we notice that the HPO trials of the last run are highly concentrated around the minimum (red line) across consecutive HPO runs, thereby confirming that previous HPO runs effectively provided meaningful information to hone in narrow regions of the hyperparameter space where accurate HydraGNN models can be found.
@@
Figure 15: Logarithm in base 10 of validation MAE for HPO trials of the first HPO run on 2,048 Frontier nodes as a function of wall-clock time expressed in seconds. The red line shows the cumulative minimum across wall-clock time.
Figure 16: Logarithm in base 10 of validation MAE for HPO trials of the second HPO run on 3,072 Frontier nodes as a function of wall-clock time expressed in seconds. The red line shows the cumulative minimum across wall-clock time.
Figure 17: Logarithm in base 10 of validation MAE for HPO trials of the third HPO run on 4,096 Frontier nodes as a function of wall-clock time expressed in seconds. The red line shows the cumulative minimum across wall-clock time.
Figure 18: Logarithm in base 10 of validation MAE for HPO trials of the fourth HPO run on 8,560 Frontier nodes as a function of wall-clock time expressed in seconds. The red line shows the cumulative minimum across wall-clock time.
@@
To characterize the dynamic resource behavior of HPO as multiple trials with varying model configurations are running in parallel, telemetry data collected by Omnistat was used to track the GPU memory usage for each trial as a function of time. Fig. 19 highlights these memory traces for all trials initiated during the final HPO exercise using 8,560 Frontier nodes during a 6-hour run. Each line on the plot corresponds to one of 390 trials initiated and we observe a dynamic high water mark peaking at 99.9% of available memory. The variability of memory utilization across different trials is due to the fact that different groups of GPUs (associated with different HPO trials) train HydraGNN models of different sizes, which affects the amount of GPU memory engaged at different stages of the model training.
@@
Figure 19: Max GPU HBM memory consumption traces sampled via Omnistat telemetry harness during final HPO exercise using 8,560 Frontier nodes (68,480 GCDs) executed on OLCF-Frontier.
@@ IV-E Energy Profiling
To quantify the energy usage as a function of different model sizes, three training epochs of the SMALL, MEDIUM, and LARGE model configurations listed in Table III were completed with the Omnistat telemetry tool sampling at one second intervals. Measurements consider the entire run duration including I/O for the initial data loading process. Each model executed using 1,024 GPUs on 128 compute nodes which is the minimum node count needed to accommodate memory requirements for the LARGE model configuration. The resulting GPU energy measurements as a function of model size are summarized in Table V. While total energy scales with the execution time, note that GPU utilization (occupancy) also influences the energy consumed. Table V includes mean utilization observed across all 1,024 GPUs. The LARGE case showed the highest GPU utilization–around 89%. The underling power histories used to compute total energy consumed for each model configuration are shown in Fig. 20. From these plots, we see evidence of the underlying training process with three epoch cycles visible in the power response. Furthermore, the increased GPU utilization for the larger models leads to increased GPU power demand with the LARGE model encountering peak power measurements in excess of 520 watts (W) (the peak TDP power for the AMD MI250 socket is 560W).
@@
Figure 20: GPU Energy use over time for three models – SMALL (top), MEDIUM (middle), and LARGE (bottom). Each line represents one AMD Instinct MI250x.
In Figure 21 we show a scatterplot of the GPU energy consumption collected using Omnistat telemetry for each HPO trial executed during the four consecutive HPO runs against the number of models in the parameters, which clearly shows a linear trend between the two quantities.
Figure 21: Energy consumption of each HPO trial of the four consecutive HPO run as a function of the number of model parameters. The red line denotes the estimated trend of a linear regression model.
We used this analysis to obtain energy-efficiency in the development and pretraining of GFMs as described in the following Section IV-F.
@@
IV-F Energy-efficient full training of best performing HydraGNN models identified by scalable HPO
To ensure trustworthiness of our GFM, we refined the pretraining of an ensemble of twenty HPO trials selected from the four consecutive HPO runs and use them for the ensemble UQ. To this effect, we sorted the HPO trials for increasing values of the validation MAE. The results showed that nine HPO trials were outperforming all the other in terms of accuracy, with a validation MAE below 0.11. The HydraGNN architectures associated with these HPO trials have been identified as worth being further trained and thus included in the ensemble for UQ as first tier. For the selection of the remaining eleven HPO trials to be included in the ensemble, we noticed that thirty-two HPO trials were clustered within a second tier with a narrow range of the validation MAE between 0.12 and 0.125. Although these models were very close to each other in terms of accuracy, they were quite different in terms of energy consumed during the HPO run. Therefore, to ensure energy efficiency, we decided to apply a second screening of the HPO trials based on energy consumption. To this end, we performed local sorting restricted only among these thirty-two HPO trials based on increasing amount of energy consumed. The eleven models with the lowest energy consumption have been selected and included in the ensemble for UQ as second tier. In Figure 22 we show the scatterplot of the energy consumption against the validation MAE for each HPO trial executed during the four consecutive HPO runs. The nine HPO trials belonging to the first tier (in red) sit in the left most part of the plot with lowest validation MAE, whereas the eleven HPO trials of the second tier (in pink) sit along the Pareto front of energy consumption vs. validation MAE. The plot also clearly shows that all the HPO trials selected (first and second tier) sit on the side of the Pareto front that partially favors preserving optimal accuracy over energy efficiency.
All twenty selected HPO trials selected use the EGNN as MPNN layer. Among the different types of MPNN layers tested, the EGNN is an equivariant model. Since equivariant models are supposed to be more data-efficient than non-equivariant models by taking full advantage of symmetries to achieve the maximal accuracy with the minimal requirement on the data, the fact that the automated HPO favors EGNN seems reasonable.